Soybean meal is 16% protein; cornmeal is 8% protein.?
How many lbs of each should be mixed together in order to get 320-lb mixture that is 13% protein.

Answer by mananth(16949)

percent ----------------quantity
soyabean 16 ----------------xLb
cornmeal 8 ----------------320-xLb
Mixture 13.00% ---------------- 320

16x+8(320-x)=13 *320

16x+2560-8x=4160
16x-8x=4160-2560
8x=1600
/8
x = 200 Lb 16.00% soyabean
120 Lb 8.00% cornmeal